OpenStudy (anonymous):

the length of a rectangle is half the width. The perimeter is 86ft. Find the width of the rectangle

OpenStudy (trisaba):

\[l =\frac{ 1 }{ 2 }w\]

OpenStudy (trisaba):

\[P =2l+2w\]

OpenStudy (anonymous):

@Vocaloid can you explain it to me?

OpenStudy (anonymous):

k?????????

OpenStudy (trisaba):

substitute l with 1/2 w

OpenStudy (trisaba):

what is the new perimeter formula

OpenStudy (trisaba):

l is L L IS Length

OpenStudy (anonymous):

1/2=86f?

OpenStudy (trisaba):

no just sub

OpenStudy (anonymous):

f+f+1/2=86f?

OpenStudy (trisaba):

where are you getting f from

OpenStudy (anonymous):

feet

OpenStudy (anonymous):

can you give the equation already please.

OpenStudy (trisaba):

\[P=86\]

OpenStudy (trisaba):

\[86=2(\frac{ 1 }{ 2 }W)+2W\]

OpenStudy (trisaba):

SIMPLIFY

OpenStudy (anonymous):

3w=86

OpenStudy (anonymous):

i'm getting 28.66

OpenStudy (anonymous):

@Trisaba ?

OpenStudy (anonymous):

@Vocaloid

OpenStudy (anonymous):

@Mehek14

OpenStudy (anonymous):

@misty1212

Vocaloid (vocaloid):

right

OpenStudy (anonymous):

no its .67 28.67

OpenStudy (anonymous):

@Vocaloid

Vocaloid (vocaloid):

right.... it depends on how you round it... 86/3 = 28.6666666 (repeated) this can be rounded to 28.67 or 28.66 depending on the rounding method used...

OpenStudy (anonymous):

ok thanks.